A Complete Bibliography of Publications in Reviews of Modern Physics (1930–1939)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

02 November 2023 Version 1.06

Title word cross-reference

γ [Hal36].

-Rays [Hal36].

1930 [CDE^{+}31].

Absolute [Cla39a, Mil33]. Absorption [Hal36]. Acoustics [Knu34]. Additions [Mul30b]. Adsorption [Bec35]. Affecting [Bla33a, Bla33b]. Allied [Swa39]. Alloys [NS38]. Altitudes [Jes39]. Analysis [AG39, Bla33a, Bla33b, Dun39]. Application [Eck30]. Architectural [Knu34]. Aspects [Bri35]. Atmosphere [SW39]. atmospheric [Hes39]. Atomic [Dun39, UT35, Wol32]. Atoms [Bro33, Dum33, Mul32, Ric37]. Attraction [Fal31]. Aurora [Hew37]. Aurorae [Hul37].

B [Bet37]. Band [Mul30a, Mul30b, Mul31, Mul32]. Barometric [CB39]. Behavior [MM39, Opp39]. Below [Wil39]. Bibliography [Gib32]. Bodies [Dum33]. Brownian [BS34]. Burst [SG39]. Bursts [FS38].

C [LB37]. Calculation [Bar38]. Cell [Gro33]. Centrifuging [Bea38]. Changes [For39]. Characteristics [Fle31]. Classification [Gib32]. Collision [CM31, Mor32]. Collisions [Bro33]. Commission [CDE^{+}31]. Complete [Bar38]. component [Joh39]. Conductivity [Wil31, Nix32]. Constants [CDE^{+}31, Dun39]. Continued [Bre33]. Continuous [Smi34]. Copper [Gro33]. Copper-Cuprous-Oxide [Gro33]. Correction [Bla33b, MW36a]. Corrections [Mul30b]. Correlation [EL35, Noy33]. Cosmic [AEM^{+}39, BUL39, BR39, Bot39, Cla39a, Cla39b, CB39, CG39, For39, FS38, Hei39, Hes39, Joh38, MM39, NA39, Opp39, SR39, VG39, Wil39, Wol39, MNS39, MC39, Val39]. Cosmic-Ray [AEM^{+}39, Bot39, Cla39a, CB39, For39, FS38, Joh38, NA39, SR39, Hes39, MNS39, MC39]. Cosmology [Rob33]. Crystal [Bla33a, Bla33b]. Crystals [FSS32, Hug36, Zwi34]. Cuprous [Gro33].

deep [AG39]. definite [CM31]. Determination [Mil33, Val39]. Developments [Knu34, Lin33]. Diary [Bra31]. Diatomic [Cra34]. Different [Cla39a]. Diffraction [Bro36, Ste30]. Dirac [HL38]. Directional [SR39]. Discharge [Loe36]. Discharges [CL30, LC31]. Discontinuous [McK30, McK31]. Discrepancy [Dun39]. Discussion [Ano39a, Ano39b, Ano39c, Ano39d, Ano39e, Opp39]. Disintegration [EL35, Ros39]. Disorder [NS38]. Dispersion [Bre32, Bre33, KB32, Lad33]. Distribution [Jon39]. Double [Bea32]. Drift [Mil33]. During [Gib32]. Dynamics [Bet37, Eck30, LB37].

Early [Gib32]. Earth [Mil33]. earthmagnetic [Hes39]. Effect [Cra34, Lin33]. Effects [Bar35, Joh38, MW36a, MW36b, MNS39]. Elastic [McM39]. Electric [Bea32]. Electrical [CL30, LC31, Wil31]. Electrically [Lad33]. Electrodynamics [Swa30]. Electrolytes [Fal31]. Electron [Bar35, Bec35, Bro36, HL38, McM39, Ric37, Mul32]. Electron-Inertia [Bar35]. Electronic [Sla34]. Electrons [Bro33, Dum33, MM39, Opp39, Smy31]. Elementary [Dar34a, Dar34b]. Elements [Gib32]. Elevations [Kor39]. Elliptic [Bar38]. Emission [Bec35, Dus30, Dus31]. Empirical [Mul31]. Energy [Jon39, MM39, Opp39, SG39, Val39]. Entropies [FSS32]. Errata [DB34b, Kem30]. Erratum [Dar34b]. Errors [DB34a, DB34b]. Ether [Mil33]. Ether-Drift [Mil33]. Evidence [Joh39]. Excess [Jon39]. Excited [Lad33]. Experiment [Mil33]. Experimental [LB37]. Experiments [BR39]. Explosion [Hei39]. Extensive [AEM^{+}39]. External [Lin33].

Factor [Kim32]. Factors [Bla33a, Bla33b]. Facts [Hew37]. Far [Ran38].

Faraday [Bra31]. field [CM31]. Fine [SR39]. Flames [Wil31]. Forbidden [Bow36]. force [CM31]. forces [Mar39]. Foreword [Ano39f]. Frequency [SG39]. Function [SG39]. Fundamental [CL30, LC31].

galvano [SF31]. galvano- [SF31]. Gas [Bro36]. Gases [Cla39a, CL30, Lad33, LC31, McM39]. General [Hof32, Kem29, Kem30, KH30]. Generalized [Bar38]. Geomagnetic [Joh38]. Ground [Wil39]. Group [Eck30, RM36]. Gyromagnetic [Bar35].

hard [Joh39]. Helium [Sat36]. Henry [Mag31]. High [Bea38, Bri35, Bro39, Jes39, Kor39, MM39, Opp39]. Hydrogen [UT35].

Ideas [Fal31]. II [Hes39, KH30, LC31, Mor32]. IIa [Mul30a, Mul30b]. IIb [Mul30a, Mul30b]. IIc [Mul31]. III [Mul32]. Inertia [Bar35]. Infra [Ran38]. Infra-Red [Ran38]. Infrared [Den31]. Inner [Ric37]. Integral [Bar38]. Intensities [Bla33a, Bla33b]. Intensity [Cla39b, CB39, For39, Joh38, SR39, VG39, Hes39, MC39]. Interesting [Bri35]. Interionic [Fal31]. International [CDE^{+}31]. Interpretation [Mul30a, Mul30b, Mul31, Mul32]. Ionization [Bro39, Cla39a, Ric37, Smy31]. Ionosphere [Mim37]. Isotope [UT35].

Joseph [Mag31].

Layers [Cla39b]. Level [Cla39a, CB39, MNS39]. Light [Déj36]. Limit [BS34]. Line [Gib32]. Linear [Dum33]. Lines [Bow36, MW36a, MW36b]. Liquid [Sat36]. Liquids [Ste30]. Low [Smy31].

Magnetic [Bea32, Hul37]. Magnetism [McK30, McK31]. Measurement [Bri39]. Measuring [BS34]. Mechanics [CM31, Dar34a, Dar34b, FSS32, Kem29, Kem30, KH30, Mor32]. Mechanism [Loe36]. Media [McK30, McK31]. Mesotron [SW39]. Mesotrons [Jon39, Ros39]. Metals [Sla34, SF31]. Meteorological [Kim32]. Methods [Bla33a, Bla33b]. Molecular [Cra34, Rod31]. Molecules [Bro36, Den31, Mul32, RM36]. Momenta [Dum33]. Monatomic [Eck30]. Motion [BS34, Mil33]. Multiple [Ric37]. Music [Fle31].

Narrow [Bot39]. National [Bri39]. Natural [BS34]. Nature [NA39, Wil39]. Night [Déj36]. Notions [Dar34a, Dar34b]. Nuclear [BB36, Bet37, EL35, LB37, Bet37, LB37]. Nuclei [BB36]. Numbers [Mul32]. Numerical [Bar38].

Observations [Gib32]. Ocean [CG39]. Optical [Con37, KB32]. Orbits [BUL39]. Order [NS38]. Order-Disorder [NS38]. Oxide [Gro33].

Pacific [CG39]. Papers [Gib32]. Part [Bec35, CL30, Den31, Gib32, Kem29, Kem30, KH30, LC31, Mor32, Mul31, Mul32, Smi34, Zwi34]. Particles [Bro39, NA39, CM31, Joh39]. Particular [FSS32]. Parts [Bre33, Mul30a, Mul30b]. Pattern [SR39]. Penetrating [Swa39]. Periodic [BUL39, VG39]. Phenomena [Bri35, LC31, Ric39, Swa39, Hes39, MC39, SF31]. Photo [Nix32]. Photo-conductivity [Nix32]. Photochemistry [Noy33]. Photoconductivity [Hug36]. Photoelectric [Gro33, Hal36, Lin33]. Physical [Fle31, Sat36]. Physics [BB36, Bet37, LB37, Mim37, Zwi34]. Polyatomic [Den31, RM36]. Positive [Jon39]. Power [Con37]. Present [Wol39]. Pressure [Bri35, CB39, MW36a, MW36b]. Pressures [FSS32]. Primary [BUL39, Joh39, Val39]. Principal [Fal31]. Principles [Kem29, Kem30, KH30]. Problem [Loe36]. Processes [BS34, CL30, CM31, EL35, Mor32, Smy31]. Produced [Swa39]. Production [SW39]. Products [Smy31]. Properties [Sat36]. protons [Joh39]. Published [Gib32].

Quantitative [Bro33]. Quantum [Bre32, Bre33, CM31, Dar34a, Dar34b, Eck30, Fer32, Kem29, Kem30, KH30, Mor32, Mul32, Smi34, VS35].

Radiation [BUL39, Cla39b, Fer32, Kim32, MM39, Opp39, VG39]. Radioactive [CDE^{+}31]. Radium [CDE^{+}31]. Radium-Standards [CDE^{+}31]. Ray [AEM^{+}39, Bla33a, Bla33b, Bot39, Cla39a, CB39, Dum33, For39, FS38, Joh38, NA39, SR39, Smi34, Ste30, Wol32, Hes39, MNS39, MC39]. Rays [BR39, CG39, Hal36, Hei39, Hes39, Rod31, Swa39, Wil39, Wol39, AG39, Val39]. Rectifier [Gro33]. Recurrence [MC39]. Red [Ran38]. Reference [FSS32]. Reflection [Bla33a, Bla33b]. Refraction [Bea32]. relation [Hes39]. Relativistic [Rob33]. Relativity [Hof32, Swa30]. Report [CDE^{+}31]. Revaluation [Dun39]. Revealed [Dum33]. Rotatory [Con37].

Scattering [Dum33, McM39, Wol32, CM31]. Sea [Cla39a, CB39, MNS39]. Seasonal [Jes39, MNS39]. Several [Bla33a, Bla33b]. Shells [Ric37]. Showers [AEM^{+}39, Bot39, FS38, Hei39, Swa39]. Sidereal [Wol39]. significance [Hes39]. Sky [Déj36]. Solar [Kim32, Wol39, Hes39]. Solid [Dum33, Sat36]. Some [BR39, Fle31]. Spark [Loe36]. Specific [Bro39]. Spectra [Cra34, Den31, Gib32, Mul30a, Mul30b, Mul31, Mul32]. Spectral [MW36a, MW36b]. Spectroscopy [Noy33, Ran38]. Spectrum [Smi34, Val39]. Speech [Fle31]. Speed [Bea38, Bro39, Smy31]. Stable [BUL39]. Standards [Bri39, CDE^{+}31]. States [BB36, Mul32]. Static [Loe36]. Stationary [BB36]. Statistical [DB34a, DB34b, FSS32, SF31]. Status [Wol39]. Strong [Fal31]. Structure [SR39, Sla34, Wol32]. Structures [Bla33a, Bla33b]. Studies [Jes39, Kor39]. Study [Bro33, Lin33]. Superconductivity [SW35]. Supersonic [Ric39]. Survey [Hew37, CL30].

Systems [Eck30, Gib32].

Temperature [CB39]. Terrestrial [Hul37]. Their [Mul32, Hes39]. Theoretical [Bet37]. Theoretically [Bri35]. Theories [Con37, Hew37]. Theory [BUL39, Bre32, Bre33, DB34a, DB34b, Fal31, Fer32, Hal36, Hei39, HL38, RM36, Smi34, VG39, VS35, Eck30, SF31]. Thermionic [Bec35, Dus30, Dus31, Bec35]. thermoelectric [SF31]. thermomagnetic [SF31]. Thick [Cla39b]. Time [Wol39, Hes39]. Transformations [NS38]. Two [UT35]. types [Mul31].

Valence [VS35]. Value [Cla39a]. Vapor [FSS32]. Variation [CB39, Wol39]. Variations [CB39, Hes39, Hul37, VG39]. VI [Bre33]. Vibrations [RM36]. VII [Bre33].

Waals [Mar39]. Weight [UT35]. Wide [For39, VG39]. World [For39, VG39]. World-Wide [For39, VG39].

X [Bla33a, Bla33b, Dum33, Hal36, Smi34, Ste30, Wol32]. X-Ray [Bla33a, Bla33b, Dum33, Smi34, Ste30, Wol32]. X-Rays [Hal36].

Zeeman [Cra34].

References

Auger:1939:ECR

[AEM^{+}39] Pierre Auger, P. Ehrenfest, R. Maze, J. Daudin, and Robley A. Fréon. Extensive cosmic-ray showers. Reviews of Modern Physics, 11(3–4):288–291, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

          org/doi/10.1103/RevModPhys.11.288;   http://rmp.aps.org/
          abstract/RMP/v11/i3-4/p288_1.

Auger:1939:ADR

[AG39] Pierre Auger and Thérèse Grivet. Analysis of deep rays. Reviews of Modern Physics, 11(3–4):232–234, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

          0756. URL http://link.aps.org/doi/10.1103/RevModPhys.11.
          232;   http://rmp.aps.org/abstract/RMP/v11/i3-4/p232_1.

Anonymous:1939:Da

[Ano39a] Anonymous. Discussion. Reviews of Modern Physics, 11(3– 4):189–190, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.11.189;   http://rmp.aps.org/
abstract/RMP/v11/i3-4/p189_1.

Anonymous:1939:Db

[Ano39b] Anonymous. Discussion. Reviews of Modern Physics, 11(3– 4):220–221, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.11.220;   http://rmp.aps.org/
abstract/RMP/v11/i3-4/p220_1.

Anonymous:1939:Dc

[Ano39c] Anonymous. Discussion. Reviews of Modern Physics, 11(3–4): 231, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.11.231;   http://rmp.aps.org/
abstract/RMP/v11/i3-4/p231_1.

Anonymous:1939:De

[Ano39d] Anonymous. Discussion. Reviews of Modern Physics, 11(3–4): 281, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.11.281;   http://rmp.aps.org/
abstract/RMP/v11/i3-4/p281_1.

Anonymous:1939:Df

[Ano39e] Anonymous. Discussion. Reviews of Modern Physics, 11(3–4): 287, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.11.287;   http://rmp.aps.org/
abstract/RMP/v11/i3-4/p287_1.

Anonymous:1939:F

[Ano39f] Anonymous. Foreword. Reviews of Modern Physics, 11(3–4): 122, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.11.122;   http://rmp.aps.org/
abstract/RMP/v11/i3-4/p122_1.

Barnett:1935:GEI

[Bar35] S. J. Barnett. Gyromagnetic and electron-inertia effects. Reviews of Modern Physics, 7(2):129–166, April 1935. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.7.129;
http://rmp.aps.org/abstract/RMP/v7/i2/p129_1.

Bartky:1938:NCG

[Bar38] Walter Bartky. Numerical calculation of a generalized complete elliptic integral. Reviews of Modern Physics, 10(4):264– 269, October 1938. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.10.264;   http://rmp.aps.org/
abstract/RMP/v10/i4/p264_1.

Bethe:1936:NPS

[BB36] H. A. Bethe and R. F. Bacher. Nuclear physics A. Stationary states of nuclei. Reviews of Modern Physics, 8(2):82–229, April 1936. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/doi/10.

1103/RevModPhys.8.82;   http://rmp.aps.org/abstract/RMP/
v8/i2/p82_1.

Beams:1932:EMD

[Bea32] J. W. Beams. Electric and magnetic double refraction. Reviews of Modern Physics, 4(1):133–172, January 1932. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.4.133;
http://rmp.aps.org/abstract/RMP/v4/i1/p133_1.

Beams:1938:HSC

[Bea38] J. W. Beams. High speed centrifuging. Reviews of Modern Physics, 10(4):245–263, October 1938. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.10.245;   http://rmp.aps.org/
abstract/RMP/v10/i4/p245_1.

Becker:1935:TEE

[Bec35] J. A. Becker. Thermionic electron emission and adsorption Part I. Thermionic emission. Reviews of Modern Physics, 7(2):95–128, April 1935. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/doi/10.

1103/RevModPhys.7.95;   http://rmp.aps.org/abstract/RMP/
v7/i2/p95_1.

Bethe:1937:NPB

[Bet37] H. A. Bethe. Nuclear physics B. Nuclear dynamics, theoretical. Reviews of Modern Physics, 9(2):69–244, April 1937. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.9.
69;   http://rmp.aps.org/abstract/RMP/v9/i2/p69_1.

Blake:1933:FARa

[Bla33a] F. C. Blake. On the factors affecting the reflection intensities by the several methods of X-ray analysis of crystal structures. Reviews of Modern Physics, 5(3):169–202, July 1933. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.5.169;
http://rmp.aps.org/abstract/RMP/v5/i3/p169_1.

Blake:1933:FARb

[Bla33b] F. C. Blake. On the factors affecting the reflection intensities by the several methods of X-ray analysis of crystal structures. A correction. Reviews of Modern Physics, 5(4):288, October 1933. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.5.
288;   http://rmp.aps.org/abstract/RMP/v5/i4/p288_1.

Bothe:1939:NCR

[Bot39] W. Bothe. Narrow cosmic-ray showers. Reviews of Modern Physics, 11(3–4):282–287, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.11.282;   http://rmp.aps.org/
abstract/RMP/v11/i3-4/p282_1.

Bowen:1936:FL

[Bow36] I. S. Bowen. Forbidden lines. Reviews of Modern Physics, 8(2):55– 81, April 1936. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/doi/10.

1103/RevModPhys.8.55;   http://rmp.aps.org/abstract/RMP/
v8/i2/p55_1.

Blackett:1939:SRE

[BR39] P. M. S. Blackett and B. Rossi. Some recent experiments on cosmic rays. Reviews of Modern Physics, 11(3–4):277– 281, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.11.277;   http://rmp.aps.org/
abstract/RMP/v11/i3-4/p277_1.

Bragg:1931:FD

[Bra31] W. H. Bragg. Faraday’s diary. Reviews of Modern Physics, 3 (4):449–463, October 1931. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.3.449;   http://rmp.
aps.org/abstract/RMP/v3/i4/p449_1.

Breit:1932:QTD

[Bre32] G. Breit. Quantum theory of dispersion. Reviews of Modern Physics, 4(3):504–576, July 1932. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.4.504;   http://rmp.
aps.org/abstract/RMP/v4/i3/p504_1.

Breit:1933:QTD

[Bre33] G. Breit. Quantum theory of dispersion (continued). Parts VI and VII. Reviews of Modern Physics, 5(2):91–140, April 1933. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.5.
91;   http://rmp.aps.org/abstract/RMP/v5/i2/p91_1.

Bridgman:1935:TIA

[Bri35] P. W. Bridgman. Theoretically interesting aspects of high pressure phenomena. Reviews of Modern Physics, 7(1):1–33, January 1935. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.7.1;   http://rmp.aps.org/abstract/RMP/v7/i1/
p1_1.

Briggs:1939:NSM

[Bri39] Lyman J. Briggs. The national standards of measurement. Reviews of Modern Physics, 11(2):111–120, April 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.11.
111;   http://rmp.aps.org/abstract/RMP/v11/i2/p111_1.

Brode:1933:QSC

[Bro33] Robert B. Brode. The quantitative study of the collisions of electrons with atoms. Reviews of Modern Physics, 5(4):257–279, October 1933. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/doi/

10.1103/RevModPhys.5.257;   http://rmp.aps.org/abstract/
RMP/v5/i4/p257_1.

Brockway:1936:EDG

[Bro36] L. O. Brockway. Electron diffraction by gas molecules. Reviews of Modern Physics, 8(3):231–266, July 1936. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.8.231;
http://rmp.aps.org/abstract/RMP/v8/i3/p231_1.

Brode:1939:SIH

[Bro39] Robert B. Brode. The specific ionization of high speed particles. Reviews of Modern Physics, 11(3–4):222–229, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.11.
222;   http://rmp.aps.org/abstract/RMP/v11/i3-4/p222_1.

Barnes:1934:BMN

[BS34] R. Bowling Barnes and S. Silverman. Brownian motion as a natural limit to all measuring processes. Reviews of Modern Physics, 6(3):162–192, July 1934. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.6.162;   http://rmp.
aps.org/abstract/RMP/v6/i3/p162_1.

Banos:1939:SPO

[BUL39] Alfredo Baños, Jr., Héctor Uribe, and Jaime Lifshitz. Stable periodic orbits in the theory of primary cosmic radiation. Reviews of Modern Physics, 11(3–4):137–148, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.11.
137;   http://rmp.aps.org/abstract/RMP/v11/i3-4/p137_1.

Clay:1939:VCR

[CB39] J. Clay and E. M. Bruins. Variations of cosmic-ray intensity with variation of barometric pressure and temperature at sea level. Reviews of Modern Physics, 11(3–4):158–159, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.11.
158;   http://rmp.aps.org/abstract/RMP/v11/i3-4/p158_1.

Curie:1931:RCR

[CDE^{+}31] M. Curie, A. Debierne, A. S. Eve, H. Geiger, O. Hahn, S. C. Lind, St. Meyer, E. Rutherford, and E. Schweidler. The radioactive constants as of 1930 Report of the International Radium-Standards Commission. Reviews of Modern Physics, 3(3):427–445, July 1931. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.3.
427;   http://rmp.aps.org/abstract/RMP/v3/i3/p427_1.

Compton:1939:CRP

[CG39] A. H. Compton and P. S. Gill. Cosmic rays on the Pacific Ocean. Reviews of Modern Physics, 11(3–4):136, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.11.
136;   http://rmp.aps.org/abstract/RMP/v11/i3-4/p136_1.

Compton:1930:EDG

[CL30] Karl T. Compton and Irving Langmuir. Electrical discharges in gases. Part I. Survey of fundamental processes. Reviews of Modern Physics, 2(2):123–242, April 1930. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.2.123;
http://rmp.aps.org/abstract/RMP/v2/i2/p123_1.

Clay:1939:AVC

[Cla39a] J. Clay. The absolute value of cosmic-ray ionization at sea level in different gases. Reviews of Modern Physics, 11(3– 4):123–127, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.11.123;   http://rmp.aps.org/
abstract/RMP/v11/i3-4/p123_1.

Clay:1939:ICR

[Cla39b] J. Clay. The intensity of cosmic radiation under thick layers. Reviews of Modern Physics, 11(3–4):128–135, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.11.
128;   http://rmp.aps.org/abstract/RMP/v11/i3-4/p128_1.

Condon:1931:QMC

[CM31] E. U. Condon and P. M. Morse. Quantum mechanics of collision processes I. Scattering of particles in a definite force field.

Reviews of Modern Physics, 3(1):43–88, January 1931. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.3.
43;   http://rmp.aps.org/abstract/RMP/v3/i1/p43_1.

Condon:1937:TOR

[Con37] E. U. Condon. Theories of optical rotatory power. Reviews of Modern Physics, 9(4):432–457, October 1937. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.9.432;
http://rmp.aps.org/abstract/RMP/v9/i4/p432_1.

Crawford:1934:ZED

[Cra34] F. H. Crawford. Zeeman effect in diatomic molecular spectra. Reviews of Modern Physics, 6(2):90–117, April 1934. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.6.
90;   http://rmp.aps.org/abstract/RMP/v6/i2/p90_1.

Darrow:1934:ENQa

[Dar34a] Karl K. Darrow. Elementary notions of quantum mechanics. Reviews of Modern Physics, 6(1):23–68, January 1934. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.6.
23;   http://rmp.aps.org/abstract/RMP/v6/i1/p23_1. See er-

ratum [Dar34b].

Darrow:1934:ENQb

[Dar34b] Karl K. Darrow. Elementary notions of quantum mechanics erratum. Reviews of Modern Physics, 6(4):281, October 1934. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.6.
281;   http://rmp.aps.org/abstract/RMP/v6/i4/p281_1. See

[Dar34a].

Deming:1934:STEa

[DB34a] W. Edwards Deming and Raymond T. Birge. On the statistical theory of errors. Reviews of Modern Physics, 6(3):119–161, July 1934. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/doi/

10.1103/RevModPhys.6.119;   http://rmp.aps.org/abstract/

RMP/v6/i3/p119_1. See errata [DB34b].

Deming:1934:STEb

[DB34b] W. Edwards Deming and Raymond T. Birge. On the statistical theory of errors errata. Reviews of Modern Physics, 6(4):281, October 1934. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/doi/

10.1103/RevModPhys.6.281;   http://rmp.aps.org/abstract/
RMP/v6/i4/p281_2. See [DB34a].

Dejardin:1936:LNS

[Déj36] Georges Déjardin. The light of the night sky. Reviews of Modern Physics, 8(1):1–21, January 1936. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.8.1;   http://rmp.
aps.org/abstract/RMP/v8/i1/p1_1.

Dennison:1931:ISP

[Den31] David M. Dennison. The infrared spectra of polyatomic molecules Part I. Reviews of Modern Physics, 3(2):280–345, April 1931. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.3.280;   http://rmp.aps.org/abstract/RMP/v3/
i2/p280_1.

Dumond:1933:LME

[Dum33] Jesse W. M. Dumond. The linear momenta of electrons in atoms and in solid bodies as revealed by X-ray scattering. Reviews of Modern Physics, 5(1):1–33, January 1933. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.5.
1;   http://rmp.aps.org/abstract/RMP/v5/i1/p1_1.

Dunnington:1939:ACR

[Dun39] Frank G. Dunnington. The atomic constants: a revaluation and an analysis of the discrepancy. Reviews of Modern Physics, 11(2):65–83, April 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/doi/

10.1103/RevModPhys.11.65;   http://rmp.aps.org/abstract/
RMP/v11/i2/p65_1.

Dushman:1930:TE

[Dus30] Saul Dushman. Thermionic emission. Reviews of Modern Physics, 2(4):381–476, October 1930. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.2.381;   http://rmp.
aps.org/abstract/RMP/v2/i4/p381_1.

Dushman:1931:TE

[Dus31] Saul Dushman. Thermionic emission. Reviews of Modern Physics, 3(1):190, January 1931. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.3.190;   http://rmp.aps.org/
abstract/RMP/v3/i1/p190_2.

Eckart:1930:AGT

[Eck30] Carl Eckart. The application of group theory to the quantum dynamics of monatomic systems. Reviews of Modern Physics, 2(3):305–380, July 1930. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/doi/

10.1103/RevModPhys.2.305;   http://rmp.aps.org/abstract/
RMP/v2/i3/p305_1.

Evans:1935:CND

[EL35] Robley D. Evans and M. Stanley Livingston. A correlation of nuclear disintegration processes. Reviews of Modern Physics, 7(3):229–236, July 1935. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/doi/

10.1103/RevModPhys.7.229;   http://rmp.aps.org/abstract/
RMP/v7/i3/p229_1.

Falkenhagen:1931:PII

[Fal31] H. Falkenhagen. The principal ideas in the interionic attraction theory of strong electrolytes. Reviews of Modern Physics, 3(3):412–426, July 1931. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/doi/

10.1103/RevModPhys.3.412;   http://rmp.aps.org/abstract/
RMP/v3/i3/p412_1.

Fermi:1932:QTR

[Fer32] Enrico Fermi. Quantum theory of radiation. Reviews of Modern Physics, 4(1):87–132, January 1, 1932. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.4.87;   http://rmp.
aps.org/abstract/RMP/v4/i1/p87_1.

Fletcher:1931:SPC

[Fle31] Harvey Fletcher. Some physical characteristics of speech and music. Reviews of Modern Physics, 3(2):258–279, April 1931. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.3.
258;   http://rmp.aps.org/abstract/RMP/v3/i2/p258_1.

Forbush:1939:WWC

[For39] S. E. Forbush. World-wide changes in cosmic-ray intensity. Reviews of Modern Physics, 11(3–4):168–172, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.11.
168;   http://rmp.aps.org/abstract/RMP/v11/i3-4/p168_1.

Froman:1938:CRS

[FS38] Darol K. Froman and J. C. Stearns. Cosmic-ray showers and bursts. Reviews of Modern Physics, 10(3):133–192, July 1938. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.10.
133;   http://rmp.aps.org/abstract/RMP/v10/i3/p133_1.

Fowler:1932:SMP

[FSS32] R. H. Fowler, F. R. S., and T. E. Sterne. Statistical mechanics with particular reference to the vapor pressures and entropies of crystals. Reviews of Modern Physics, 4(4):635–722, October 1932. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.4.
635;   http://rmp.aps.org/abstract/RMP/v4/i4/p635_1.

Gibbs:1932:LSE

[Gib32] R. C. Gibbs. Line spectra of the elements Part I. Early observations and systems of classification bibliography I. Papers published during 1920–1931. Reviews of Modern Physics, 4(2):278–470, April 1932. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/doi/

10.1103/RevModPhys.4.278;   http://rmp.aps.org/abstract/
RMP/v4/i2/p278_1.

Grondahl:1933:CCO

[Gro33] L. O. Grondahl. The copper-cuprous-oxide rectifier and photoelectric cell. Reviews of Modern Physics, 5(2):141–168, April 1933.

CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.5.141;   http://rmp.aps.org/abstract/RMP/v5/
i2/p141_1.

Hall:1936:TPA

[Hal36] Harvey Hall. The theory of photoelectric absorption for X-rays and γ-rays. Reviews of Modern Physics, 8(4):358–397, October 1936. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/doi/

10.1103/RevModPhys.8.358;   http://rmp.aps.org/abstract/
RMP/v8/i4/p358_1.

Heisenberg:1939:TES

[Hei39] W. Heisenberg. On the theory of explosion showers in cosmic rays. Reviews of Modern Physics, 11(3–4):241, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.11.
241;   http://rmp.aps.org/abstract/RMP/v11/i3-4/p241_1.

Hess:1939:ITV

[Hes39] Victor F. Hess. II. Time variations of cosmic rays. The significance of variations in cosmic-ray intensity and their relation to solar, earthmagnetic and atmospheric phenomena. Reviews of Modern Physics, 11(3–4):153–157, July 1939. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.11.
153;   http://rmp.aps.org/abstract/RMP/v11/i3-4/p153_1.

Hewson:1937:SFT

[Hew37] E. W. Hewson. A survey of the facts and the theories of the aurora. Reviews of Modern Physics, 9(4):403–431, October 1937. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.9.
403;   http://rmp.aps.org/abstract/RMP/v9/i4/p403_1.

Hill:1938:DET

[HL38] E. L. Hill and R. Landshoff. The Dirac electron theory. Reviews of Modern Physics, 10(2):87–132, April 1938. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.10.87;
http://rmp.aps.org/abstract/RMP/v10/i2/p87_1.

Hoffmann:1932:GR

[Hof32] Banesh Hoffmann. On general relativity. Reviews of Modern Physics, 4(1):173–204, January 1932. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.4.173;   http://rmp.
aps.org/abstract/RMP/v4/i1/p173_1.

Hughes:1936:PC

[Hug36] A. L. Hughes. Photoconductivity in crystals. Reviews of Modern Physics, 8(3):294–315, July 1936. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.8.294;   http://rmp.
aps.org/abstract/RMP/v8/i3/p294_1.

Hulburt:1937:TMV

[Hul37] E. O. Hulburt. Terrestrial magnetic variations and aurorae. Reviews of Modern Physics, 9(1):44–68, January 1937. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.9.
44;   http://rmp.aps.org/abstract/RMP/v9/i1/p44_1.

Jesse:1939:SSH

[Jes39] William P. Jesse. Seasonal studies at high altitudes. Reviews of Modern Physics, 11(3–4):167, July 1939. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.11.
167;   http://rmp.aps.org/abstract/RMP/v11/i3-4/p167_1.

Johnson:1938:CRI

[Joh38] Thomas H. Johnson. Cosmic-ray intensity and geomagnetic effects. Reviews of Modern Physics, 10(4):193–244, October 1938. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.10.
193;   http://rmp.aps.org/abstract/RMP/v10/i4/p193_1.

Johnson:1939:EPP

[Joh39] Thomas H. Johnson. Evidence that protons are the primary particles of the hard component. Reviews of Modern Physics, 11 (3–4):208–210, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.11.208;   http://rmp.aps.org/
abstract/RMP/v11/i3-4/p208_1.

Jones:1939:EDP

[Jon39] Haydn Jones. Energy distribution and positive excess of mesotrons. Reviews of Modern Physics, 11(3–4):235–238, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.11.
235;   http://rmp.aps.org/abstract/RMP/v11/i3-4/p235_1.

Korff:1932:OD

[KB32] S. A. Korff and G. Breit. Optical dispersion. Reviews of Modern Physics, 4(3):471–503, July 1932. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.4.471;   http://rmp.
aps.org/abstract/RMP/v4/i3/p471_1.

Kemble:1929:GPQ

[Kem29] Edwin C. Kemble. The general principles of quantum mechanics. Part I. Reviews of Modern Physics, 1(2):157–215, October 1929. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/doi/

10.1103/RevModPhys.1.157;   http://rmp.aps.org/abstract/

RMP/v1/i2/p157_1. See errata [Kem30].

Kemble:1930:EPG

[Kem30] Edwin C. Kemble. Errata to Part I, general principles of quantum mechanics. Reviews of Modern Physics, 2(1):59, January 1930. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.2.59;   http://rmp.aps.org/abstract/RMP/

v2/i1/p59_1. See [Kem29].

Kemble:1930:GPQ

[KH30] E. C. Kemble and E. L. Hill. The general principles of quantum mechanics. Part II. Reviews of Modern Physics, 2(1):1–59, January 1930. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.2.1;   http://rmp.aps.org/abstract/RMP/v2/i1/
p1_1.

Kimball:1932:SRM

[Kim32] Herbert H. Kimball. Solar radiation as a meteorological factor. Reviews of Modern Physics, 4(2):259–277, April 1932. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.4.259;
http://rmp.aps.org/abstract/RMP/v4/i2/p259_1.

Knudsen:1934:RDA

[Knu34] Vern O. Knudsen. Recent developments in architectural acoustics. Reviews of Modern Physics, 6(1):1–22, January 1934. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.6.
1;   http://rmp.aps.org/abstract/RMP/v6/i1/p1_1.

Korff:1939:RSH

[Kor39] S. A. Korff. Recent studies at high elevations. Reviews of Modern Physics, 11(3–4):211–219, July 1939. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.11.
211;   http://rmp.aps.org/abstract/RMP/v11/i3-4/p211_1.

Ladenburg:1933:DEE

[Lad33] Rudolf Ladenburg. Dispersion in electrically excited gases. Reviews of Modern Physics, 5(4):243–256, October 1933. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.5.
243;   http://rmp.aps.org/abstract/RMP/v5/i4/p243_1.

Livingston:1937:NPC

[LB37] M. Stanley Livingston and H. A. Bethe. Nuclear physics C. Nuclear dynamics, experimental. Reviews of Modern Physics, 9(3):245–390, July 1937. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/doi/

10.1103/RevModPhys.9.245;   http://rmp.aps.org/abstract/
RMP/v9/i3/p245_1.

Langmuir:1931:EDG

[LC31] Irving Langmuir and Karl T. Compton. Electrical discharges in gases Part II. Fundamental phenomena in electrical discharges. Reviews of Modern Physics, 3(2):191–257, April 1931. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.3.191;
http://rmp.aps.org/abstract/RMP/v3/i2/p191_1.

Linford:1933:RDS

[Lin33] Leon B. Linford. Recent developments in the study of the external photoelectric effect. Reviews of Modern Physics, 5(1):34–61, January 1933. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/doi/10.

1103/RevModPhys.5.34;   http://rmp.aps.org/abstract/RMP/
v5/i1/p34_1.

Loeb:1936:PMS

[Loe36] Leonard B. Loeb. The problem of the mechanism of static spark discharge. Reviews of Modern Physics, 8(3):267–293, July 1936. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.8.267;   http://rmp.aps.org/abstract/RMP/v8/
i3/p267_1.

Magie:1931:JH

[Mag31] W. F. Magie. Joseph Henry. Reviews of Modern Physics, 3 (4):465–495, October 1931. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.3.465;   http://rmp.
aps.org/abstract/RMP/v3/i4/p465_1.

Margenau:1939:VWF

[Mar39] H. Margenau. Van der Waals forces. Reviews of Modern Physics, 11(1):1–35, January 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.11.1;   http://rmp.
aps.org/abstract/RMP/v11/i1/p1_1.

Monk:1939:RPC

[MC39] A. T. Monk and A. H. Compton. Recurrence phenomena in cosmic-ray intensity. Reviews of Modern Physics, 11(3–4): 173–179, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.11.173;   http://rmp.aps.org/
abstract/RMP/v11/i3-4/p173_1.

McKeehan:1930:MDM

[McK30] L. W. McKeehan. Magnetism in discontinuous media. Reviews of Modern Physics, 2(4):477–505, October 1930. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.2.477;
http://rmp.aps.org/abstract/RMP/v2/i4/p477_1.

McKeehan:1931:MDM

[McK31] L. W. McKeehan. Magnetism in discontinuous media. Reviews of Modern Physics, 3(1):190, January 1931. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.3.190;
http://rmp.aps.org/abstract/RMP/v3/i1/p190_1.

McMillen:1939:EES

[McM39] J. Howard McMillen. Elastic electron scattering in gases. Reviews of Modern Physics, 11(2):84–110, April 1939. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.11.84;
http://rmp.aps.org/abstract/RMP/v11/i2/p84_1.

Miller:1933:EDE

[Mil33] Dayton C. Miller. The ether-drift experiment and the determination of the absolute motion of the Earth. Reviews of Modern Physics, 5(3):203–242, July 1933. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.5.203;
http://rmp.aps.org/abstract/RMP/v5/i3/p203_1.

Mimno:1937:PI

[Mim37] Harry Rowe Mimno. The physics of the ionosphere. Reviews of Modern Physics, 9(1):1–43, January 1937. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.9.
1;   http://rmp.aps.org/abstract/RMP/v9/i1/p1_1.

Montgomery:1939:BHE

[MM39] C. G. Montgomery and D. D. Montgomery. The behavior of high energy electrons in the cosmic radiation. Reviews of Modern Physics, 11(3–4):255–264, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.11.255;   http://rmp.aps.org/
abstract/RMP/v11/i3-4/p255_1. See discussion [Opp39].

Millikan:1939:SCR

[MNS39] R. A. Millikan, H. V. Neher, and D. O. Smith. Seasonal cosmic-ray effects at sea level. Reviews of Modern Physics, 11 (3–4):166, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.11.166;   http://rmp.aps.org/
abstract/RMP/v11/i3-4/p166_1.

Morse:1932:QMC

[Mor32] Philip M. Morse. Quantum mechanics of collision processes Part II. Reviews of Modern Physics, 4(3):577–634, July 1932. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.4.
577;   http://rmp.aps.org/abstract/RMP/v4/i3/p577_1.

Mulliken:1930:IBSa

[Mul30a] Robert S. Mulliken. The interpretation of band spectra. Parts I, IIa, IIb. Reviews of Modern Physics, 2(1):60–115, January 1930. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.2.60;   http://rmp.aps.org/abstract/RMP/
v2/i1/p60_1.

Mulliken:1930:IBSb

[Mul30b] Robert S. Mulliken. The interpretation of band spectra, Parts I, IIa, IIb. Additions and corrections. Reviews of Modern Physics, 2(4):506–508, October 1930. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.2.506;   http://rmp.
aps.org/abstract/RMP/v2/i4/p506_1.

Mulliken:1931:IBS

[Mul31] Robert S. Mulliken. The interpretation of band spectra. Part IIc. Empirical band types. Reviews of Modern Physics, 3(1):89–155, January 1931. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/doi/10.

1103/RevModPhys.3.89;   http://rmp.aps.org/abstract/RMP/
v3/i1/p89_1.

Mulliken:1932:IBS

[Mul32] Robert S. Mulliken. The interpretation of band spectra Part III. Electron quantum numbers and states of molecules and their atoms. Reviews of Modern Physics, 4(1):1–86, January 1932. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.4.
1;   http://rmp.aps.org/abstract/RMP/v4/i1/p1_1.

Margenau:1936:CPE

[MW36a] H. Margenau and W. W. Watson. A correction: Pressure effects on spectral lines. Reviews of Modern Physics, 8(4):398, October 1936. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/doi/

10.1103/RevModPhys.8.398;   http://rmp.aps.org/abstract/
RMP/v8/i4/p398_1.

Margenau:1936:PES

[MW36b] Henry Margenau and William W. Watson. Pressure effects on spectral lines. Reviews of Modern Physics, 8(1):22–53, January 1936. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/doi/10.

1103/RevModPhys.8.22;   http://rmp.aps.org/abstract/RMP/
v8/i1/p22_1.

Neddermeyer:1939:NCR

[NA39] Seth H. Neddermeyer and Carl D. Anderson. Nature of cosmic-ray particles. Reviews of Modern Physics, 11(3–4):191– 207, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.11.191;   http://rmp.aps.org/
abstract/RMP/v11/i3-4/p191_1.

Nix:1932:PC

[Nix32] Foster C. Nix. Photo-conductivity. Reviews of Modern Physics, 4(4):723–766, October 1932. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.4.723;   http://rmp.
aps.org/abstract/RMP/v4/i4/p723_1.

Noyes:1933:CSP

[Noy33] W. Albert Noyes, Jr. The correlation of spectroscopy and photochemistry. Reviews of Modern Physics, 5(4):280–287, October 1933. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/doi/

10.1103/RevModPhys.5.280;   http://rmp.aps.org/abstract/
RMP/v5/i4/p280_1.

Nix:1938:ODT

[NS38] Foster C. Nix and William Shockley. Order-disorder transformations in alloys. Reviews of Modern Physics, 10(1):1–71, January 1938. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/doi/10.

1103/RevModPhys.10.1;   http://rmp.aps.org/abstract/RMP/
v10/i1/p1_1.

Oppenheimer:1939:DBH

[Opp39] J. R. Oppenheimer. Discussion [the behavior of high energy electrons in the cosmic radiation]. Reviews of Modern Physics, 11 (3–4):264–266, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.11.264;   http://rmp.aps.org/
abstract/RMP/v11/i3-4/p264_1. See [MM39].

Randall:1938:SFI

[Ran38] H. M. Randall. The spectroscopy of the far infra-red. Reviews of Modern Physics, 10(1):72–85, January 1938. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.10.72;
http://rmp.aps.org/abstract/RMP/v10/i1/p72_1.

Richtmyer:1937:MII

[Ric37] F. K. Richtmyer. The multiple ionization of inner electron shells of atoms. Reviews of Modern Physics, 9(4):391–402, October 1937. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.9.391;   http://rmp.aps.org/abstract/RMP/v9/
i4/p391_1.

Richards:1939:SP

[Ric39] William T. Richards. Supersonic phenomena. Reviews of Modern Physics, 11(1):36–64, January 1939. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.11.36;
http://rmp.aps.org/abstract/RMP/v11/i1/p36_1.

Rosenthal:1936:GTV

[RM36] Jenny E. Rosenthal and G. M. Murphy. Group theory and the vibrations of polyatomic molecules. Reviews of Modern Physics, 8(4):317–346, October 1936. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.8.317;   http://rmp.
aps.org/abstract/RMP/v8/i4/p317_1.

Robertson:1933:RC

[Rob33] H. P. Robertson. Relativistic cosmology. Reviews of Modern Physics, 5(1):62–90, January 1933. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.5.62;   http://rmp.
aps.org/abstract/RMP/v5/i1/p62_1.

Rodebush:1931:MR

[Rod31] Worth H. Rodebush. Molecular rays. Reviews of Modern Physics, 3(3):392–411, July 1931. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.3.392;   http://rmp.aps.org/
abstract/RMP/v3/i3/p392_1.

Rossi:1939:DM

[Ros39] Bruno Rossi. The disintegration of mesotrons. Reviews of Modern Physics, 11(3–4):296–303, July 1939. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.11.
296;   http://rmp.aps.org/abstract/RMP/v11/i3-4/p296_1.

Satterly:1936:PPS

[Sat36] John Satterly. The physical properties of solid and liquid helium. Reviews of Modern Physics, 8(4):347–357, October 1936. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.8.
347;   http://rmp.aps.org/abstract/RMP/v8/i4/p347_1.

Sommerfeld:1931:STT

[SF31] A. Sommerfeld and N. H. Frank. The statistical theory of thermoelectric, galvano- and thermomagnetic phenomena in metals. Reviews of Modern Physics, 3(1):1–42, January 1931. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.3.
1;   http://rmp.aps.org/abstract/RMP/v3/i1/p1_1.

Schein:1939:BFF

[SG39] Marcel Schein and Piara S. Gill. Burst frequency as a function of energy. Reviews of Modern Physics, 11(3–4):267–276, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.11.267;   http://rmp.aps.org/
abstract/RMP/v11/i3-4/p267_1.

Slater:1934:ESM

[Sla34] J. C. Slater. The electronic structure of metals. Reviews of Modern Physics, 6(4):209–280, October 1934. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.6.209;
http://rmp.aps.org/abstract/RMP/v6/i4/p209_1.

Smith:1934:QTC

[Smi34] Lloyd P. Smith. Quantum theory of the continuous X-ray spectrum. Part I. Reviews of Modern Physics, 6(2):69–89, April 1934. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.6.
69;   http://rmp.aps.org/abstract/RMP/v6/i2/p69_1.

Smyth:1931:PPI

[Smy31] H. D. Smyth. Products and processes of ionization by low speed electrons. Reviews of Modern Physics, 3(3):347–391, July 1931. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.3.
347;   http://rmp.aps.org/abstract/RMP/v3/i3/p347_1.

Schremp:1939:FSP

[SR39] E. J. Schremp and H. S. Ribner. On the fine structure pattern of directional cosmic-ray intensity. Reviews of Modern Physics, 11 (3–4):149–152, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.11.149;   http://rmp.aps.org/
abstract/RMP/v11/i3-4/p149_1.

Stewart:1930:XRD

[Ste30] G. W. Stewart. X-ray diffraction in liquids. Reviews of Modern Physics, 2(1):116–122, January 1930. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.2.116;
http://rmp.aps.org/abstract/RMP/v2/i1/p116_1.

Smith:1935:S

[SW35] H. Grayson Smith and J. O. Wilhelm. Superconductivity. Reviews of Modern Physics, 7(4):237–271, October 1935. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.7.
237;   http://rmp.aps.org/abstract/RMP/v7/i4/p237_1.

Schein:1939:MPA

[SW39] Marcel Schein and Volney C. Wilson. Mesotron production in the atmosphere. Reviews of Modern Physics, 11(3–4):292– 295, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.11.292;   http://rmp.aps.org/
abstract/RMP/v11/i3-4/p292_1.

Swann:1930:RE

[Swa30] W. F. G. Swann. Relativity and electrodynamics. Reviews of Modern Physics, 2(3):243–304, July 1930. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.2.243;
http://rmp.aps.org/abstract/RMP/v2/i3/p243_1.

Swann:1939:SPP

[Swa39] W. F. G. Swann. Showers produced by penetrating rays and allied phenomena. Reviews of Modern Physics, 11(3–4):242– 254, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.11.242;   http://rmp.aps.org/
abstract/RMP/v11/i3-4/p242_1.

Urey:1935:HIA

[UT35] Harold C. Urey and Gordon K. Teal. The hydrogen isotope of atomic weight two. Reviews of Modern Physics, 7(1):34–94, January 1935. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/doi/10.

1103/RevModPhys.7.34;   http://rmp.aps.org/abstract/RMP/
v7/i1/p34_1.

Vallarta:1939:DES

[Val39] M. S. Vallarta. The determination of the energy spectrum of primary cosmic rays. Reviews of Modern Physics, 11(3– 4):239–240, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.11.239;   http://rmp.aps.org/
abstract/RMP/v11/i3-4/p239_1.

Vallarta:1939:TWW

[VG39] M. S. Vallarta and O. Godart. A theory of world-wide periodic variations of the intensity of cosmic radiation. Reviews of Modern Physics, 11(3–4):180–189, July 1939. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.11.
180;   http://rmp.aps.org/abstract/RMP/v11/i3-4/p180_1.

VanVleck:1935:QTV

[VS35] J. H. Van Vleck and Albert Sherman. The quantum theory of valence. Reviews of Modern Physics, 7(3):167–228, July 1935. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.7.
167;   http://rmp.aps.org/abstract/RMP/v7/i3/p167_1.

Wilson:1931:ECF

[Wil31] H. A. Wilson. Electrical conductivity of flames. Reviews of Modern Physics, 3(1):156–189, January 1931. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.3.156;
http://rmp.aps.org/abstract/RMP/v3/i1/p156_1.

Wilson:1939:NCR

[Wil39] Volney C. Wilson. Nature of cosmic rays below ground. Reviews of Modern Physics, 11(3–4):230–231, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.11.
230;   http://rmp.aps.org/abstract/RMP/v11/i3-4/p230_1.

Wollan:1932:XRS

[Wol32] E. O. Wollan. X-ray scattering and atomic structure. Reviews of Modern Physics, 4(2):205–258, April 1932. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.4.205;
http://rmp.aps.org/abstract/RMP/v4/i2/p205_1.

Wollan:1939:PSS

[Wol39] E. O. Wollan. Present status of solar and sidereal time variation of cosmic rays. Reviews of Modern Physics, 11(3–4): 160–165, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.11.160;   http://rmp.aps.org/
abstract/RMP/v11/i3-4/p160_1.

Zwicky:1934:PCP

[Zwi34] F. Zwicky. On the physics of crystals. Part I. Reviews of Modern Physics, 6(3):193–208, July 1934. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.6.193;
http://rmp.aps.org/abstract/RMP/v6/i3/p193_1.